Starting from rest, a boulder rolls down a hill with constant acceleration and travels 2.00 during the first second.
Answer #1
s = v+u /2 *t
u = 0 since it is the first second.
2 = v /2 * 1
v = 4 m/s
Answer #2
x = vt + 0.5at^2

So for the 1st second:

x = 2.00m
t= 1s
v = 0 (this is initial velocity)

2 = 0.5a
a = 2/0.5
a=4m/s^2

Acceleration is constant so we can use it to solve the questions

A. Use the same equation as before, except:
t=2s
a=4m/s/s
v=0
x=?

x= 0.5(4)*(2)^2
= 8 m

So it traveled 8 meters all together, 6 meters from t=1s to t=2sec ( x(1)=2 and x(2)=8 )
Answer #3

as
a = 2.00m/sec^2
t = 1.00sec
v = a*t = 2.00*1.00 = 2 m/sec (speed)
i hope it helps you :)